# Distinguishing radiation mechanisms and particle populations in blazar jets through long-term multi-band monitoring with RINGO3 and Fermi

### v1

@@ 3.3 Optical spectral index – gamma-ray
Fig. 4 shows the optical spectral index vs $\gamma$-ray flux correlations for each object in our sample, and Table 7 in the appendix shows their correlation strengths and significances. The $\gamma$-ray flux data is independent of the optical, meaning the optical spectral index could be calculated using all three optical flux colours (b*g*r*), avoiding correlating errors (see Appendix A for more details). The error calculation for the spectral index per epoch, and the determination of linear/logarithmic fit preference was the same as detailed in Section 3.2. To reduce the chance of false correlations, we remove repeated instances where the same $\gamma$-ray flux value has been assigned to multiple optical epochs.
Of the 28 sources, 15 show significant correlations between the optical spectral index and $\gamma$-ray flux. The strengths of these correlations range from very weak ($\pm$0.164) to very strong ($\pm$0.897) with both positive and negative correlation strengths.
Of the 15 BL Lac types, eight showed significant behaviour (2 HSPs, 1 ISP, and all 5 LSPs). Additionally, all but two of the significantly correlated sources were negative in strength, implying an increase in $\gamma$-ray emission correlated with a decreasing optical spectral index (i.e. the sources became optically bluer when $\gamma$-ray brighter). Furthermore, five objects demonstrated the BSWB trend. Two sources, namely A0 0235+164 (LSP) and OJ287 (LSP), showed the opposite trend; these objects became optically redder when brighter in $\gamma$-rays, but followed the linear relationship. For the 13 FSRQs, seven showed correlated behaviour between the optical spectral index and $\gamma$-ray flux. When brighter in $\gamma$-rays one of these correlations showed the BSWB trend, while the remaining six showed the RSWB trend.
Figure 4: As Fig. 3 but for spectral index vs $\gamma$-ray flux. The Spearman rank coefficients associated with each correlation are given in Table 7.
@@ 3.4 Optical – gamma-ray time lags
The Discrete Correlation Function (DCF; Edelson & Krolik, 1988) was used to explore time lags between the optical and $\gamma$-ray data as described in McCall et al. (2024). The DCF points are fitted using the GaussianProcessRegressor module and the Rational Quadratic kernel from scikit-learn in Python (Pedregosa et al., 2011). The maximum absolute value from this fit is chosen as the lag. The error on this lag is the average cadence of the optical light curve. The optical r*-band flux is shifted with respect to the $\gamma$-ray flux, meaning a negative lag implies the $\gamma$-ray emission is leading the optical, and a positive lag implies the optical emission is leading the $\gamma$-ray.
The results are shown in Table 8 in the appendix. If a lag is inconsistent with zero at the $3\sigma$ level and has a strength coefficient of greater than 0.40 (see Table 2) we label the lag as potentially significant. The sources that meet these criteria are 1ES 1959+650, ON231, PKS 1510-089, and PKS B1622-297. To determine whether the measured lags were real, they were applied to the optical data and the correlation was re-tested using the Spearman rank coefficient. We make the assumption that if the lags were real, the optical and $\gamma$-ray fluxes would become, or remain, significantly correlated with a larger strength value. These results are shown in Table 3 where the Spearman rank correlation statistics before and after applying the detected lags are shown. We found that the correlations for PKS 1510-089 and PKS B1622-297 did not increase in strength after shifting, whereas for both 1ES 1959+650 and ON231, the optical–$\gamma$-ray correlations became significantly correlated. The coefficients suggested an inverse correlation for 1ES 1959+650, and direct for ON231.
Source Lag $\Delta$Lag p c
@@ 4 Discussion
Table 4: Results of the different correlations explored in this work. The source name and subclass are given, along with ‘Yes’ or ‘No’ to describe whether or not the given correlation was statistically significant.
@@ 4.1 Optical–gamma-ray analysis
Sources with high synchrotron peaks showed fewer significant correlations (33 per cent for HSPs). The number of $\gamma$-ray–optical correlations increases to 67 per cent of ISP sources and 89 per cent of LSP objects (which includes all BL Lac objects and 85 per cent of FSRQs). Jermak et al. (2016) explored optical-$\gamma$-ray correlations for a sample of 15 blazars monitored with the RINGO2 and DIPOL polarimeters and determined that significant positive correlations were found in 68 per cent of cases. When considering object sub-classes, they found 43 per cent, 50 per cent, 88 per cent, and 60 per cent of HBL, IBL, LBL, and FSRQ sources, respectively, showed positively correlated behaviour. The overall results are therefore similar between the RINGO2 and RINGO3 analyses. The decrease in strength of correlations for blazar subclass (based on location of synchrotron peak) and optical-$\gamma$-ray emission can be understood as host-galaxy contamination in the case of HSPs, which may dilute the optical behaviour of the jet and as such result in weaker correlations with the $\gamma$-ray flux (Gaur, 2014; Chang et al., 2019; Otero-Santos et al., 2022). However, one cannot rule out the possibility that the intrinsic properties of HSPs differ from that of ISP and LSP objects, leading to weaker correlations.
@@
Figures 1 and 2 show the distribution of sources in the optical and $\gamma$-flux phase space as a function of apparent and absolute magnitude, respectively. Based on the optical and $\gamma$-ray flux properties, most sources in this sample occupied one of two regions in these plots. These regions are attributed to the FSRQ/BL Lac classification of blazars; with FSRQs typically displaying brighter $\gamma$-ray fluxes than BL Lacs. There are three sources, however, that appear to occupy the region in between these two subclasses of blazar; A0 0235+164, S4 0954+64 and 4C 09.57. A0 0235+164 was originally classified as a BL Lac LSP source by Spinrad & Smith (1975) due to its featureless spectrum, however, Ruan et al. (2014) indicated that A0 0235+164 may belong to the FSRQ transitional class of blazars. This supports the conclusions of Ackermann et al. (2012) who model the A0 0235+164 SED during a flaring episode in 2008-2009 and find that the source’s isotropic luminosity is more indicative of FSRQs than BL Lacs, and the dominant mechanism for the high energy emission is more likely to be external Compton processes; a signature of FSRQs, rather than synchrotron self-Compton. Similar to A0 0235+164, S4 0954+65 was also classified as a BL Lac object (Stickel et al., 1991). Ghisellini et al. (2011) classify this source as a low-frequency peaked blazar (LBL) due to the absence of prominent emission lines and the appearance of its SED. However, Hervet et al. (2016) classify the sources as an FSRQ by analysing the kinematic features of its radio jet. Furthermore, MAGIC Collaboration et al. (2018) model the multiwavelength emission of S4 0954+65 and compare it to other sources, concluding that it shows many behavioural similarities to FSRQs and other suspected transitional/masquerading BL Lac objects. Ghisellini et al. (2011) classify 4C 09.57 as a BL Lac LSP object based on the shape of its SED, however, they also observe emission lines with equivalent widths up to 12.5Å. 4C 09.57 may also be a transitional object between the two subclasses (Uemura et al., 2017).
@@
The correlations presented in Figs. 1 and 2 also agree with the results of Hovatta et al. (2014) in that the flux-flux correlations appear tighter for HSP and ISP sources compared to that of LSP objects (BL Lacs and FSRQs). In the SSC case, the low- and high-energy emission originates from the same region of the blazar jet so the emission is subject to the same level of Doppler boosting. This would mean those sources which have $\gamma$-ray production dominated by SSC emission should show tight correlations between the optical and $\gamma$-ray fluxes. Any EC emission should be more strongly boosted, obscuring any linear dependence between the low- and high-energy emission (Dermer, 1995). In this EC case, the flux-flux correlations would appear more scattered. Our results support HSP and ISP sources having SSC as the dominant $\gamma$-ray emission mechanism, whereas LSP sources show significant EC emission.
4.2 Spectral analysis (vs optical and gamma-ray)
@@
Fig. 3 shows the spectral index vs optical flux diagrams for those sources with optical and $\gamma$-ray data. Significant correlations were found for 28 of the 30 sources (15 BL Lacs and all 13 FSRQs). For the BL Lac objects, all 15 showed BWB behaviour with six (40 per cent) displaying a logarithmic trend, indicating the stabilisation of the colour at higher fluxes (bluer-stable-when-brighter; BSWB). The FSRQs consisted of two linear RWB and BWB relationships while the rest (85 per cent) showed stable trends. One of the stable trends was BSWB, while the remaining eight were RSWB. The Spearman rank strengths of all significant correlations ranged from ±$\pm 0.195$±$\pm 0.953$ (weak–very strong).
@@
This same analysis was used to look for non-linear relationships between the optical spectral index and $\gamma$-ray flux, the results of which are shown in Fig. 4 with best fits and correlation statistics shown in Table 7. We found 15 sources to have significant correlations (54 per cent), eight BL Lacs (two HSP, one ISP, and five LSP) and seven FRSQs. Only three of the BL Lac objects displayed linear relationships (one BWB and two RWB), while the remaining five BL Lacs and all seven FSRQs showed stable features (all BSWB for the BL Lacs types and one FSRQ, and six RSWB for the FSRQs).
@@ 4.3 Time lag analysis
Our analysis differs from that of related work in that we look for long-term lag behaviour between the optical and $\gamma$-ray bands, rather than individual flares. However, our analysis is in agreement with recent work (Cohen et al., 2014; Liodakis et al., 2019; de Jaeger et al., 2023) in that we find little evidence for temporal lags between the optical and $\gamma$-ray bands that are not consistent with zero days at the $3\sigma$ level. This means our data is suggestive of SSC processes dominating the jet flux, but further analysis into the characteristics of individual optical flares, and any gamma-ray correlations, could reveal more detail.
@@ 4.4 Correlations summary
13 sources show significant correlations between the spectral index and optical flux but not the spectral index and $\gamma$-ray flux. Of these 13, six (three HSPs, one ISP, and two FSRQs) do not show optical and $\gamma$-ray flux correlations, possibly due to relatively inactive or indiscernible $\gamma$-ray behaviour during our observations. The remaining seven sources (one HSP, one ISP, one LSP, and four FSRQs) do show significant optical and $\gamma$-ray flux correlations with no detected time lag. This implies their $\gamma$-ray brightening episodes are optically colourless; a possible indication of $\gamma$-ray emission originating from seed photon fields outside the jet and a signature of EC processes.
@@
Two objects (1ES 1959+650; HSP, and A0 0235+164; LSP) demonstrate significant spectral index vs $\gamma$-ray flux correlations, but not spectral index vs optical flux correlations. Furthermore, 1ES 1959+650 shows no significant correlation between the optical and $\gamma$-ray fluxes whereas A0 0235+164 does. In the case of 1ES 1959+650, the correlations indicate that optical colour variability from jet emission may be present, but would not be temporally consistent with any optical flux, but rather consistent with preceding or delayed $\gamma$-ray activity. Our time lag result for this source hinted at such behaviour but was ultimately discouraged due to unevenly sampled optical data. More observations of this object with regular sampling need to be obtained, ideally during a range of activity states, in order to make more definitive conclusions. For A0 0235+164, the majority of our observations occurred during a prolonged optical and $\gamma$-ray heightened state. It is possible that our optical observations only detected the flat region of a logarithmic RSWB colour trend and therefore appeared colourless, as not enough data was obtained during quiescent and/or low states. Furthermore, it has been previously reported that this object has a much brighter accretion disc than other LSP BL Lac types, and is more comparable to that of FSRQs (Ghisellini et al., 2010; Zhang et al., 2022). This would explain why RWB spectral index vs $\gamma$-ray flux behaviour was observed, in contrast to the generally accepted, almost exclusively BWB behaviour of BL Lac type blazars.
Finally, one object (OJ287; LSP) showed significantly correlated optical and $\gamma$-ray fluxes, but conflicting trends in the spectral index vs optical flux (BSWB) and spectral index vs $\gamma$-ray flux (RWB). We note that spectral index vs $\gamma$-ray flux is very weakly correlated, and may be slightly skewed, attributed to an expected accretion disk impact flare by a hypothesised secondary SMBH companion (Lehto & Valtonen, 1996). This will be explored in more detail in a future publication utilising the full RINGO3 photopolarimetric data set.
@@ 5 Conclusions
In this work we have distinguished between radiation mechanisms and particle populations in blazars jets using a large photometric dataset of blazars. RINGO3 is also a polarimeter and polarimetric observations of blazars can be instrumental in disentangling the jet’s synchrotron emission from any thermal contributions from other parts of the AGN. These polarimetric properties will be explored in the next publication. The results presented in this work encourage further high-cadence photopolarimetric observations of all blazar subclasses to ensure adequate monitoring of all activity states.
@@ Acknowledgements
The Liverpool Telescope is operated on the island of La Palma by Liverpool John Moores University in the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofisica de Canarias with financial support from the UKRI Science and Technology Facilities Council (STFC) (ST/T00147X/1).
IA acknowledges financial support from the grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033 to the Instituto de Astrofísica de Andalucía-CSIC and from MICIN grants PID2019-107847RB-C44 and PID2022-139117NB-C44.
GPL is supported by Royal Society Dorothy Hodgkin Fellowship (grant Nos. DHF-R1-221175 and DHFERE-221005)
UBA acknowledges the receipt of a FAPERJ State Scientist Fellowship nr. E-26/200.532/2023 and a CNPq Productivity Fellowship nr. 309053/2022-6
@@
All data presented here are available through the Liverpool Telescope data archive at https://telescope.livjm.ac.uk/. The data underlying this article will be shared upon reasonable request to the corresponding author.

### v2

@@ 3.3 Optical spectral index – gamma-ray
Fig. 4 shows the optical spectral index vs. $\gamma$-ray flux correlations for each object in our sample, and Table 7 in the appendix shows their correlation strengths and significances. The $\gamma$-ray flux data are independent of the optical, meaning the optical spectral index could be calculated using all three optical flux colours (b*g*r*), avoiding correlating errors (see Appendix A for more details). The error calculation for the spectral index per epoch, and the determination of linear/logarithmic fit preference was the same as detailed in Section 3.2. To reduce the chance of false correlations, we remove repeated instances where the same $\gamma$-ray flux value has been assigned to multiple optical epochs.
[…]
Of the 15 BL Lac type sources, eight showed significant behaviour (2 HSPs, 1 ISP, and all 5 LSPs). Additionally, all but two of the significantly correlated sources were negative in strength, implying an increase in $\gamma$-ray emission correlated with a decreasing optical spectral index (i.e. the sources became optically bluer when $\gamma$-ray brighter). Furthermore, five objects demonstrated the BSWB trend. Two sources, namely A0 0235+164 (LSP) and OJ287 (LSP), showed the opposite trend; these objects became optically redder when brighter in $\gamma$-rays, but followed the linear relationship. For the 13 FSRQs, seven showed correlated behaviour between the optical spectral index and $\gamma$-ray flux. When brighter in $\gamma$-rays one of these correlations showed the BSWB trend, while the remaining six showed the RSWB trend.
Figure 4: As Fig. 3 but for spectral index vs. $\gamma$-ray flux. The Spearman rank coefficients associated with each correlation are given in Table 7.
@@ 3.4 Optical – gamma-ray time lags
The Discrete Correlation Function (DCF; Edelson & Krolik, 1988) was used to explore time lags between the optical and $\gamma$-ray data as described in McCall et al. (2024). The DCF points are fitted using the GaussianProcessRegressor module and the Rational Quadratic kernel from scikit-learn in Python (Pedregosa et al., 2011). The maximum absolute value from this fit is chosen as the lag. The optical r*-band flux is shifted with respect to the $\gamma$-ray flux, meaning a negative lag implies the $\gamma$-ray emission is leading the optical, and a positive lag implies the optical emission is leading the $\gamma$-ray.
The results are shown in Table 8 in the appendix. If a lag is inconsistent with zero at the $3\sigma$ level and has a strength coefficient of greater than 0.40 (see Table 2) we label the lag as potentially significant (yes*). The $\Delta$Peak column shows the error on the peak value, given as the average cadence between successive optical observations. In all cases, this value was larger than the error associated with the calculation of the DCF. The sources that meet these criteria are 1ES 1959+650, ON231, PKS 1510-089, and PKS B1622-297. To determine whether the measured lags were real, they were applied to the optical data and the correlation was re-tested using the Spearman rank coefficient. We make the assumption that if the lags were real, the optical and $\gamma$-ray fluxes would become, or remain, significantly correlated with a larger strength value. These results are shown in Table 3 where the Spearman rank correlation statistics before and after applying the detected lags are shown. We found that the correlations for PKS 1510-089 and PKS B1622-297 did not increase in strength after shifting, whereas for both 1ES 1959+650 and ON231, the optical–$\gamma$-ray correlations became significantly correlated. The coefficients suggested an inverse correlation for 1ES 1959+650, and direct for ON231.
Source Lag [days] $\Delta$Lag [days] $p$ $c$
@@ 4 Discussion
Table 4: Results of the different correlations explored in this work. The source name and subclass are given, along with ‘Yes’ or ‘No’ to describe whether or not the given correlation was statistically significant. The asterisk in the $3\sigma$ lag column indicates the potential significance of the detected lag, warranting further analysis.
@@ 4.1 Optical–gamma-ray analysis
Sources with high synchrotron peaks showed fewer significant correlations (33 per cent for HSPs). The number of $\gamma$-ray–optical correlations increases to 67 per cent of ISP sources and 89 per cent of LSP objects (which includes all BL Lac objects and 85 per cent of FSRQs). Jermak et al. (2016) explored optical-$\gamma$-ray correlations for a sample of 15 blazars monitored with the RINGO2 and DIPOL polarimeters and determined that significant positive correlations were found in 68 per cent of cases. When considering object sub-classes, they found 43 per cent, 50 per cent, 88 per cent, and 60 per cent of HBL, IBL, LBL, and FSRQ sources, respectively, showed positively correlated behaviour. The overall results are therefore similar between the RINGO2 and RINGO3 analyses. The decrease in the strength of optical–$\gamma$-ray emission correlations on increasing synchrotron peak frequency (blazar subclass) can be understood as host-galaxy contamination in the case of HSPs. The host-galaxy emission would dilute the optical behaviour of the jet and as such result in weaker correlations with the $\gamma$-ray flux (Gaur, 2014; Chang et al., 2019; Otero-Santos et al., 2022). However, one cannot rule out the possibility that the intrinsic properties of HSPs differ from those of ISP and LSP objects, leading to weaker correlations.
@@
Based on the optical and $\gamma$-ray flux properties, most sources in this sample occupied one of two regions in figures 1 and 2. These regions are attributed to the FSRQ/BL Lac classification of blazars; with FSRQs typically displaying brighter $\gamma$-ray fluxes than BL Lacs. There are three sources, however, that appear to occupy the region in between these two blazar subclasses: A0 0235+164, S4 0954+64 and 4C 09.57. A0 0235+164 was originally classified as a BL Lac LSP source by Spinrad & Smith (1975) due to its featureless spectrum, however, Ruan et al. (2014) indicated that A0 0235+164 may belong to the FSRQ transitional class of blazars. This supports the conclusions of Ackermann et al. (2012) who model the A0 0235+164 SED during a flaring episode in 2008-2009 and find that the source’s isotropic luminosity is more indicative of FSRQs than BL Lacs, and the dominant mechanism for the high energy emission is more likely to be external Compton processes; a signature of FSRQs, rather than synchrotron self-Compton. Similar to A0 0235+164, S4 0954+65 was also classified as a BL Lac object (Stickel et al., 1991). Ghisellini et al. (2011) classify this source as a low-frequency peaked blazar (LBL) due to the absence of prominent emission lines and the appearance of its SED. However, Hervet et al. (2016) classify the sources as an FSRQ by analysing the kinematic features of its radio jet. Furthermore, MAGIC Collaboration et al. (2018) model the multiwavelength emission of S4 0954+65 and compare it to other sources, concluding that it shows many behavioural similarities to FSRQs and other suspected transitional/masquerading BL Lac objects. Ghisellini et al. (2011) classify 4C 09.57 as a BL Lac LSP object based on the shape of its SED, however, they also observe emission lines with equivalent widths up to 12.5Å. 4C 09.57 may also be a transitional object between the two subclasses (Uemura et al., 2017).
@@
The correlations presented in Figs. 1 and 2 also agree with the results of Hovatta et al. (2014) in that the flux-flux correlations appear tighter for HSP and ISP sources compared to that of LSP objects (BL Lacs and FSRQs). In the SSC case, the low- and high-energy emission originates from the same region of the blazar jet so the emission is subject to the same level of Doppler boosting. This would mean those sources which have $\gamma$-ray production dominated by SSC emission should show tight correlations between the optical and $\gamma$-ray fluxes. Any EC emission should be more strongly boosted, obscuring any linear dependence between the low- and high-energy emission (Dermer, 1995). In this EC case, the flux-flux correlations would appear more scattered. Our results support HSP and ISP sources having SSC as the dominant $\gamma$-ray emission mechanism, whereas LSP sources would show significant EC emission.
4.2 Spectral analysis (vs. optical and gamma-ray)
@@
Fig. 3 shows the spectral index vs. optical flux diagrams for all sources in the sample. Significant correlations were found for 28 of the 30 sources (15 BL Lacs and all 13 FSRQs). For the BL Lac objects, all 15 showed BWB behaviour with six (40 per cent) displaying a logarithmic trend, indicating the stabilisation of the colour at higher fluxes (bluer-stable-when-brighter; BSWB). Of the FSRQs, two showed linear RWB trends and one showed a linear BWB relationship. The rest (85 per cent) showed stable trends. One of the stable trends was BSWB, while the remaining eight were RSWB. The Spearman rank strengths of all significant correlations ranged from $\sim\,|0.195|$$\sim\,|0.953|$ (weak–very strong).
@@
This same analysis was used to look for non-linear relationships between the optical spectral index and $\gamma$-ray flux, the results of which are shown in Fig. 4 with best fits and correlation statistics shown in Table 7. We found 15 sources to have significant correlations (54 per cent), eight BL Lacs (two HSP, one ISP, and five LSP) and seven FRSQs. Only three of the BL Lac objects displayed linear relationships (one BWB and two RWB), while the remaining five BL Lacs and all seven FSRQs showed stable features (all BSWB for the BL Lacs types and one FSRQ, and six RSWB for the remaining FSRQs).
@@ 4.3 Time lag analysis
Our analysis differs from that of related work in that we look for long-term lag behaviour between the optical and $\gamma$-ray bands, rather than individual flares. However, our analysis is in agreement with recent work (Cohen et al., 2014; Liodakis et al., 2019; de Jaeger et al., 2023) in that we find little evidence for temporal lags between the optical and $\gamma$-ray bands that are not consistent with zero days at the $3\sigma$ level. This means our data are suggestive of SSC processes dominating the jet flux, but further analysis into the characteristics of individual optical flares, and any gamma-ray correlations, could reveal more detail.
@@ 4.4 Correlations summary
13 sources show significant correlations between the spectral index and optical flux but not between the spectral index and $\gamma$-ray flux. Of these 13, six (three HSPs, one ISP, and two FSRQs) do not show optical and $\gamma$-ray flux correlations, possibly due to relatively inactive or indiscernible $\gamma$-ray behaviour during our observations. The remaining seven sources (one HSP, one ISP, one LSP, and four FSRQs) do show significant optical and $\gamma$-ray flux correlations with no detected time lag. This implies their $\gamma$-ray brightening episodes are optically colourless; a possible indication of $\gamma$-ray emission originating from seed photon fields outside the jet and a signature of EC processes.
@@
Two objects (1ES 1959+650; HSP, and A0 0235+164; LSP) demonstrate significant spectral index vs. $\gamma$-ray flux correlations, but not spectral index vs. optical flux correlations. Furthermore, 1ES 1959+650 shows no significant correlation between the optical and $\gamma$-ray fluxes whereas A0 0235+164 does. In the case of 1ES 1959+650, the correlations indicate that optical colour variability from jet emission may be present, but would not be temporally consistent with any optical flux, but rather consistent with preceding or delayed $\gamma$-ray activity. Our time lag result for this source hinted at such behaviour but was ultimately disregarded due to unevenly sampled optical data. More observations of this object with regular sampling need to be obtained, ideally during a range of activity states, in order to make more definitive conclusions. For A0 0235+164, the majority of our observations occurred during a prolonged optical and $\gamma$-ray heightened state. It is possible that our optical observations only detected the flat region of a logarithmic RSWB colour trend and therefore appeared colourless, as not enough data were obtained during quiescent and/or low states. Furthermore, it has been previously reported that this object has a much brighter accretion disc than other LSP BL Lacs, and is more comparable to that of FSRQs (Ghisellini et al., 2010; Zhang et al., 2022). This would explain why RWB spectral index vs. $\gamma$-ray flux behaviour was observed, in contrast to the generally accepted, almost exclusively BWB behaviour of BL Lac type blazars.
Finally, one object (OJ287; LSP) showed significantly correlated optical and $\gamma$-ray fluxes, but conflicting trends in the spectral index vs. optical flux (BSWB) and spectral index vs. $\gamma$-ray flux (RWB). We note that spectral index vs. $\gamma$-ray flux is very weakly correlated, and may be slightly skewed, attributed to an expected accretion disk impact flare by a hypothesised secondary SMBH companion (Lehto & Valtonen, 1996). This will be explored in more detail in a future publication utilising the full RINGO3 photopolarimetric data set.
@@ 5 Conclusions
In this work, we have distinguished between radiation mechanisms and particle populations in blazars jets using a large photometric dataset of blazars. RINGO3 is also a polarimeter and polarimetric observations of blazars can be instrumental in disentangling the jet’s synchrotron emission from any thermal contributions from other parts of the AGN. These polarimetric properties will be explored in the next publication. The results presented in this work encourage further high-cadence photopolarimetric observations of all blazar subclasses to ensure adequate monitoring of all activity states.
@@ Acknowledgements
We thank the anonymous reviewer for their constructive comments. The Liverpool Telescope is operated on the island of La Palma by Liverpool John Moores University in the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofisica de Canarias with financial support from the UKRI Science and Technology Facilities Council (STFC) (ST/T00147X/1).
I.A. acknowledges financial support from the grant CEX2021-001131-S funded by MCIN/AEI/10.13039/501100011033 to the Instituto de Astrofísica de Andalucía-CSIC and from MICIN grants PID2019-107847RB-C44 and PID2022-139117NB-C44.
U.B.A. acknowledges the receipt of a FAPERJ State Scientist Fellowship nr. E-26/200.532/2023 and a CNPq Productivity Fellowship nr. 309053/2022-6
T.H. was supported by the Research Council of Finland projects 317383, 320085, 322535, and 345899.
G.P.L. is supported by Royal Society Dorothy Hodgkin Fellowship (grant Nos. DHF-R1-221175 and DHFERE-221005)
@@
All data presented here are available through the Liverpool Telescope data archive at https://telescope.astro.ljmu.ac.uk. The data underlying this article will be shared upon reasonable request to the corresponding author.